write the equation of the line passing through each of the given points with the indicated slope. (rewrite in standard from) Give your results in slope-intercept form, where possible.
(0,5), m= -3/5

Found 2 solutions by stanbon, checkley75:
Answer by stanbon(75887)

Give your results in slope-intercept form, where possible.
(0,5), m= -3/5
-------------------
The form is y = mx + b
You are told the slope is -3/5 and the b is 5
EQUATION:
y = (-3/5)x + 5
=============
Cheers,
Stan H.

Answer by checkley75(3666)

Y=mX+b WHERE m=SLOPE & b=Y INTERCEPT.
NOW REPLACE THE X & Y TERMS BY (0,5) & m BY -3/5 & SOLVE FOR b
5=-3/5*0+b
5=b WHICH IS THE Y INTERCEPT.
THUS THE EQUATION IS:
Y=-3X/5+5